# Fast, Modular, and Differentiable Framework for Machine Learning-Enhanced Molecular Simulations

### v1

@@ III.1 Classical Force Fields
In classical molecular mechanics modelling of a system, one distinguishes between non-bonded and covalently-bonded atoms, where the former interactions are composed of pair-wise potentials of van der Waals type and electrostatic Coulomb interactions. The interaction between bonded atoms, in addition to a pair-wise type potential between atoms, is modeled by introducing higher-order interactions. Some force-field parameterization additionally include corrections which are based on tabular definitions of interaction energies between four atoms, called energy correction map (CMAP) [27]. We implement the energy components needed to reproduce the most common parameterizations of AMBER [23] class force fields, and some additional components often used in CHARMM type parameterizations [27]. Fig. 4(d) schematically displays typical energy components of classical force fields, which we have implemented in DIMOS and will present in more details in the following.
III.1.1 Bonded Interactions
@@
The interaction between two atoms $i$ and $j$ that are covalently bonded is modeled by a simple harmonic potential
$U_{\mathrm{bonded}}(i,j)=k_{ij}(r_{ij}-r_{ij}^{0})^{2}$ (1)
@@
where $r_{ij}=|\vec{r}_{ij}|$ is the Euclidean distance between the atoms, $k_{ij}$ models the interaction strength and $r_{ij}^{0}$ is the equilibrium or target distance. Usually, $k_{ij}$ and $r_{ij}^{0}$ depend on the atom types involved in the bond and not explicitly on the particular atom, but we still use this more general notation for simplicity.
@@
$U_{\mathrm{angle}}(i,j,l)=k_{ijl}(\theta_{ijl}-\theta_{ijl}^{0})^{2}$ (2)
where using the unit vector $\hat{r}_{ij}=\vec{r}_{ij}/|\vec{r}_{ij}|$ one defines $\theta_{ijl}=\arccos(\hat{r}_{ij}\cdot\hat{r}_{lj})$ as the angle between the three atoms, and the other parameters are defined in analogy to the harmonic bonded potential. For some force-fields, the three-body interactions are described by the Urey-Bradley potential, which is a combination of above angle and bond potentials.
Interactions between four sequentially bonded atoms $i,j,l,m$ can be described by a torsion potential
$U_{\mathrm{torsion}}(i,j,l,m)=\sum_{n}k_{ijlm}^{n}\left[1+\cos(n\omega_{ijlm}-\omega_{ijlm}^{0})\right]$ (3)
where $\omega_{ijlm}$ is the dihedral angle between the two planes spanned by $i,j,l$ and $j,l,m$ and $n$ specifies the multiplicity. The other parameters are defined in analogy to the previous two contributions. For the torsion parameters, one distinguishes between proper and improper dihedral angles, differing whether the involved atoms are in a chain or not. Alternatively, some force fields simply model the improper dihedrals harmonically.
@@
III.1.2 Nonbonded Interactions
In a usual simulation, the most time consuming part of the calculations is related to the non-bonded atoms of the system. A priori, all atoms interact with each other, leading to $O(N^{2})$ complexity for the computations. For methods using synchronous dynamics, such as MD simulations, one routinely relies on heuristic cut-offs, where for non-integrable interactions one performs the computation both in real and reciprocal space, leading to an improved convergence [28].
@@ III.1.2 Nonbonded Interactions
The introduction of a cut-off allows one to use specialized data structures to store pairs of atoms that are within the interaction radius of each other, so-called neighborlists (more generally called edge lists in the ML literature). In the following, we will discuss the components of neighborlist and their implementation in DIMOS. After this, the two main non-bonded energy contributions of classical force fields are discussed.
Neighborlists
@@
In DIMOS, we employ a combination of Verlet neighborlist [29] and cell lists [30] to construct the Verlet lists. Neighborlists are a data structure in which one simply stores all pairs of atoms that are within a distance $r_{\mathrm{nl}}=r_{c}+r_{g}$, where $r_{c}$ is the cutoff distance of the potential, and $r_{g}$ is the so-called ghost radius [29] (set to $0.25r_{c}$ by default in DIMOS). This implies that the list contains more atom pairs than strictly needed for the computation of the interactions and one wastes some computation when calculating the distance $r$ for all pairs with $r_{\mathrm{nl}}>r>r_{c}$. However, this is more than compensated by being able to reuse the neighborlist more than once during the simulation, because each atom only moves a relatively small distance at each iteration. Due to the introduction of the ghost radius $r_{g}$, the pair list only looses its validity whenever a single atom has moved a distance $d$ greater than half the ghost radius $d>r_{g}/2$ since construction of the list. Only then, the neighborlist needs to be reconstructed. This reduces the complexity to $O(N)$ per synchronous update of all atoms whenever the neighborlist is valid, in contrast to an $O(N^{2})$ complexity otherwise.
@@
However, a naive re-construction of the neighborlist still has $O(N^{2})$ complexity, ultimately not improving the complexity of the simulation. This is because the distance between all pairs of atoms needs to be calculated to decide which ones are below $r_{\mathrm{nl}}$. An alluring solution to this problem is the partitioning of the simulation box into cells based on the spatial coordinates of the atoms. Each cell is chosen to be of at least $r_{\mathrm{nl}}$ length, and each atom is assigned to one of the cells. Then, to find all atom pairs within $r_{\mathrm{nl}}$, one simply needs to check the neighboring cells (including off-diagonal neighbors) and store all combinations of atoms within the neighborlist. The overall complexity of the algorithm is $O(N)$.
@@
By default, the end-user of DIMOS intending to run a simulation does not need to explicitly consider the use of neighborlists, since these are automatically been utilized and kept up-to-date. For some applications, as for example training an MLIP on some pre-existing data, one can obtain the atom pairs as
@@
In case the boundary conditions are not periodic, the system is, for the sake of neighborlist computation, embedded into a non-periodic box. For this new (virtual) box, the usual routines are called without wrapping around the coordinates of the cells as would be otherwise necessary with periodic boundaries.
Van der Waals and Coulomb interactions
One of the most used functions to model van der Waals interactions is the 121212126666 Lennard Jones potential
@@
The parameters $\epsilon_{ij}$ and $\sigma_{ij}$ are determined from atom specific parameters $\epsilon_{i}$ and $\sigma_{i}$ by combining rules as
@@
$S(x)=1-6x^{5}+15x^{4}-10x^{3},$ (6)
to Eq. (4), where $x=(r_{ij}-r_{s})/(r_{c}-r_{s})$ and $r_{s}<r_{c}$ is the switching distance. For simulations at constant temperature and pressure in the NPT ensemble, we have also implemented long-range dispersion corrections that approximate the contributions beyond $r_{c}$ [31].
@@
A less heuristic method to introduce a cutoff in electrostatic modelling is Ewald summation [16], for which one writes
@@
$U$ =${}_{\mathrm{dir.}}=\frac{1}{2}\sum_{i,j}\sum_{\vec{n}\in\bf n}q_{i}q_{j}\frac{\operatorname{erfc}(\alpha|\vec{r}+\vec{n}|)}{|\vec{r}+\vec{n}|}$ (10)
$U$ =${}_{\mathrm{recip.}}=\frac{1}{2\pi V}\sum_{i,j}\sum_{\vec{k}\in{\bf k}\neq 0}\frac{\exp(-(\pi\vec{k}/\alpha)^{2}+2\pi\vec{k}\cdot\vec{r}_{ij})}{\vec{k}^{2}}$ (11)
$U$ =${}_{\mathrm{self}}=-\frac{\alpha}{\sqrt{\pi}}\sum_{i}q_{i}^{2}.$ (12)
The potential is thus split into a real space contribution $U_{\mathrm{dir.}}$, a reciprocal space contribution $U_{\mathrm{recip.}}$, and the self energy $U_{\mathrm{self}}$. The sums $\sum_{\vec{n}\in\bf n}$ and $\sum_{\vec{k}\in{\bf k}\neq 0}$ run over all periodic copies of the system, respectively, all reciprocal vectors. $V$ is the volume of the box, and $\alpha$ controls the shift between $U_{\mathrm{recip.}}$ and $U_{\mathrm{self}}$.
By construction, the sums of the individual energy contributions in real and reciprocal space converge significantly faster than directly evaluating the sum. Hence, it becomes possible to choose a much smaller cutoff for both sums to reach a target accuracy, significantly reducing the number of necessary computations.
For an ideal choice of cutoff of the real and reciprocal summation as well as crossover parameter $\alpha$, it can even be shown that this reduces the computational effort from $O(N^{2})$ to $O(N^{3/2})$ per step. Note that for any choice of $\alpha$ above formulas are correct; hence, the choice of $\alpha$ only influences which cutoff $r_{c}$ and $k_{c}$ can be chosen for a desired accuracy. In practice, $\alpha$ is by default heuristically determined from the prescribed tolerance $\delta$ and a fixed cutoff $r_{c}$ (typically the one used for the van der Waals interactions) as [26]
@@
To further speed-up computation, it is possible to carry out the reciprocal sum using a fast Fourier transform (FFT) by distributing the charges onto a lattice via spline interpolation. This significantly speeds up the calculation of $U_{\mathrm{recip.}}$, and the overall complexity also reduces to $O(N\log N)$. Following OpenMM [26], we implement smooth PME [15]. For this, we also implement B-splines [33] based on the recursive definition of the splines, resulting in custom forward and backward implementation within PyTorch. The number of nodes in the mesh is chosen heuristically as $n_{\mathrm{mesh}}=2\alpha d/3\delta^{1/5}$, where $\alpha$ is again determined from Eq. (13).
Exclusions and Exceptions
@@
Atoms that are directly bonded or indirectly bonded to each other are excluded from the non-bonded energy terms. In addition, atoms connected via three bonds (such as in torsion contributions) are often treated differently. The most common way to handle these $1$-$4$ exceptions is to scale the strength of the non-bonded interactions by some common factor $<1$. Alternatively, some parameterizations of force-fields scale each exclusion by an individual factor. In DIMOS, exclusions and exceptions are handled at the level of neighborlists, where excluded interaction pairs are filtered out. Since all possible $1$-$4$ exceptions are known at simulation setup, they are handled explicitly by the non-bonded interaction method.
III.2 Machine Learning Based Descriptions of Intermolecular Interactions
@@
MLIPs do not typically introduce an explicit distinction between non-bonded and bonded atoms, and instead often build higher-order interactions between all atoms within a cutoff distance using various approaches, see Ref. [25] and references therein. Usually, these approaches include physical features in some capacity, for example a large body of work only considers features that are invariant to rotations and translations of the system and uses rather standard ML architectures with learnable parameters [34]. Another line of work explicitly constructs the architectures such that their internal features or their output transform equivariantly to the input, i.e., for example rotating the system also leads to a predictable rotation of the output [17, 35, 36].
@@
We implement the interface to two current state-of-the-art MLIP models: MACE [17] and ORB [35]. We want, however, to stress that DIMOS makes no explicit assumption on the model and is rather agnostic to the details, as long as the models are implemented in PyTorch. In Fig. 4(b) we schematically demonstrate the basic setup of message-passing graph neural network MLIPs, in which many-body interaction terms are built using successive message passing iterations. However, since we do not implement any novel MLIP architecture in DIMOS, we refer to the literature on details as well as advantages/disadvantages of the respective approaches [37, 25].
@@
Using DIMOS, the best performance is achieved if the model expects a combination of the neighborlist, distance vectors or distances, and/or shifts, since they are calculated anyway in DIMOS and recalculating (some of them) within the model leads to wasted calculations. By having the full simulation framework implemented in PyTorch, it is possible to compile parts or the whole simulation. This also extends to compiling MLIPs using torch.compile, which is currently not possible when using LAMMPS [38] or OpenMM [26] to interface the MLIP. Setting up a system modeled by a (foundational) MACE model can be achieved in a single line:
@@
In addition, there are some optional parameters, which, however, are not necessary in general. This highlights the simplicity of using DIMOS to run fast and scalable simulations using an MLIP, compared to other established approaches requiring a vast number of steps.
@@
MACE is currently one of the most prominent MLIPs in the literature [17], with diverse applications in different domains. We specifically employ the MACE material project foundation model (mace-mp[18], which is available in the small, medium, and large variant depending on the chosen parameterization. In DIMOS, we support the acceleration of MACE using NVIDIA cuEquivariance  [^4] See https://github.com/NVIDIA/cuEquivariance for more details., and activate it in all our benchmarks.
III.3 Machine Learning/Molecular Mechanics

### v2

@@ II.1 Classical Force Fields
In classical molecular mechanics modelling (MM) based on classical force fields, one distinguishes between non-bonded and covalently-bonded atoms, see Fig. 1(d), where the former interactions are commonly composed of pair-wise potentials of van der Waals type and electrostatic Coulomb interactions. The interaction between bonded atoms, in addition to a pair-wise bond potential between atoms, is modeled by introducing higher-order interactions. Some force-field parameterization additionally include corrections which are based on tabular definitions of interaction energies between four atoms, called energy correction maps (CMAP).[21]
We implement the energy components needed to reproduce the most common parameterizations of AMBER[18] class force fields, and some additional components often used in CHARMM type parameterizations.[21]
II.1.1 Bonded Interactions
@@
The interaction between two atoms $i$ and $j$ that are covalently bonded is modeled by a simple harmonic pair-wise bond potential
$U_{\mathrm{bonded}}(i,j)=k_{ij}(r_{ij}-r_{ij}^{0})^{2},$ (1)
@@
where $r_{ij}=|\vec{r}_{ij}|$ is the Euclidean distance between the atoms, $k_{ij}$ models the interaction strength and $r_{ij}^{0}$ is the equilibrium distance. Usually, $k_{ij}$ and $r_{ij}^{0}$ depend on the atom types involved in the bond and not explicitly on the particular atom, but we still use this more general notation for simplicity.
@@
$U_{\mathrm{angle}}(i,j,l)=k_{ijl}(\theta_{ijl}-\theta_{ijl}^{0})^{2},$ (2)
where using the unit vector $\hat{r}_{ij}=\vec{r}_{ij}/|\vec{r}_{ij}|$ one defines $\theta_{ijl}=\arccos(\hat{r}_{ij}\cdot\hat{r}_{lj})$ as the angle between the three atoms, and the other parameters are defined in analogy to the harmonic bonded potential. For some force-fields, the three-body interactions are described by the Urey-Bradley potential, which is a combination of the angle and bond potential.
Interactions between four sequentially bonded atoms $i,j,l,m$ are described by a torsion potential
$U_{\mathrm{torsion}}(i,j,l,m)=\sum_{n}k_{ijlm}^{n}\left[1+\cos(n\omega_{ijlm}-\omega_{ijlm}^{0})\right],$ (3)
where $\omega_{ijlm}$ is the dihedral angle between the two planes spanned by $i,j,l$ and $j,l,m$ and $n$ specifies the multiplicity. The other parameters are defined in analogy to the previous two contributions. For the torsion parameters, one distinguishes between proper and improper dihedral angles, differing whether the involved atoms are in a sequential linear chain or not. Alternatively, some force fields simply model the improper dihedrals harmonically.
@@
II.1.2 Nonbonded Interactions
@@
In many simulations, the most time consuming part of the calculations is related to the non-bonded interactions. A priori, all atoms interact with each other, leading to $O(N^{2})$ complexity for the computation of energy or forces. The introduction of a cut-off allows one to use specialized data structures to store pairs of atoms that are within the interaction radius of each other, so-called neighborlists (typically called edge lists in the ML literature). Then, the complexity of the energy or force computation becomes $O(N)$ or $O(N\log N)$, depending on the selected approximation scheme. In the following, we will discuss the components of neighborlists and their implementation in DIMOS.
Neighborlists:
@@
In DIMOS, we employ a combination of Verlet neighborlist[22] and cell lists.[23] Neighborlists are data structures in which one simply stores all pairs of atoms that are within a distance $r_{\mathrm{nl}}=r_{c}+r_{g}$, where $r_{c}$ is the cutoff distance of the potential, and $r_{g}$ is the so-called ghost radius[22] (set to $0.25r_{c}$ by default in DIMOS). This implies that the list contains more atom pairs than strictly needed for the computation of the interactions and one wastes some computation when calculating the distance $r$ for all pairs with $r_{\mathrm{nl}}>r>r_{c}$. However, this is more than compensated by being able to reuse the neighborlist more than once during the simulation. This is because each atom only moves a relatively small distance at each iteration. Due to the introduction of the ghost radius $r_{g}$, the pair list only looses its validity whenever a single atom has moved a distance $d$ greater than half the ghost radius $d>r_{g}/2$ since construction of the list. Only then, the neighborlist needs to be reconstructed. This reduces the complexity to $O(N)$ per synchronous update of all atoms whenever the neighborlist is valid, in contrast to an $O(N^{2})$ complexity otherwise.
@@
However, a naive re-construction of the neighborlist still has $O(N^{2})$ complexity, ultimately not improving the complexity of the simulation. This is because the distance between all pairs of atoms needs to be calculated to decide which ones are smaller than $r_{\mathrm{nl}}$. An alluring solution to this problem is the partitioning of the simulation box into cells based on the spatial coordinates of the atoms. Each cell is chosen to be of at least $r_{\mathrm{nl}}$ length, and each atom is assigned to one of the cells. Then, to find all atom pairs within $r_{\mathrm{nl}}$, one simply needs to check the neighboring cells (including off-diagonal neighbors) and store all combinations of atoms within the neighborlist. The overall complexity of the algorithm is $O(N)$.
@@
By default, the end-user of DIMOS does not need to explicitly consider the use of neighborlists, since these are automatically been utilized and kept up-to-date. For some applications, as for example training an MLIP on some pre-existing data, one can obtain the explicit list of atom pairs as shown in the following code snippet:
@@
In case the boundary conditions are not periodic, the system is, for the sake of neighborlist computation, embedded into a non-periodic box. For these new (virtual) boxes, the usual routines are called without wrapping around the coordinates of the cells as would be otherwise necessary with periodic boundaries.
Van der Waals interactions:
One of the most used functions to model van der Waals interactions is the 121266 Lennard Jones potential
@@
The parameters $\epsilon_{ij}$ and $\sigma_{ij}$ are determined from atom specific parameters $\epsilon_{i}$ and $\sigma_{i}$ by the Lorentz-Berthelot combining rules[24, 25] as
@@
Alternatively, also fully geometric combining can be specified.
@@
$S(x)=1-6x^{5}+15x^{4}-10x^{3}$ (6)
to Eq. (4), where $x=(r_{ij}-r_{s})/(r_{c}-r_{s})$ and $r_{s}<r_{c}$ is the switching distance. For simulations at constant temperature and pressure, we have also implemented long-range dispersion corrections that approximate the contributions beyond $r_{c}$.[26]
Coulomb interactions:
@@
A less heuristic method to introduce a cutoff in electrostatic modelling is Ewald summation,[28] for which one writes
@@
$U$ $_{\mathrm{dir.}}=\frac{1}{2}\sum_{i,j}\sum_{\vec{n}\in\bf n}q_{i}q_{j}\frac{\operatorname{erfc}(\alpha|\vec{r}+\vec{n}|)}{|\vec{r}+\vec{n}|}$ (10)
$U$ $_{\mathrm{recip.}}=\frac{1}{2\pi V}\sum_{i,j}\sum_{\vec{k}\in{\bf k}\neq 0}\frac{\exp(-(\pi\vec{k}/\alpha)^{2}+2\pi\vec{k}\cdot\vec{r}_{ij})}{\vec{k}^{2}}$ (11)
$U$ $_{\mathrm{self}}=-\frac{\alpha}{\sqrt{\pi}}\sum_{i}q_{i}^{2}.$ (12)
The potential is thus split into a real space contribution $U_{\mathrm{dir.}}$, a reciprocal space contribution $U_{\mathrm{recip.}}$, and the self energy $U_{\mathrm{self}}$. The direct and reciprocal space sums run over all periodic copies of the system, respectively, all reciprocal vectors. $V$ is the volume of the simulation box, and $\alpha$ controls the shift between $U_{\mathrm{recip.}}$ and $U_{\mathrm{self}}$.
By construction, the sums of the individual energy contributions in real and reciprocal space converge significantly faster than directly evaluating the sum. Hence, it becomes possible to introduce truncations of both sums, in practice introducing a cut-off $r_{c}$ in real, respectively, $k_{c}$ in reciprocal space. Compared to the naive sum, a much smaller cutoff for both sums can be used to reach a target accuracy, significantly reducing the number of necessary computations.
For an ideal choice of cutoff of the real and reciprocal summation as well as crossover parameter $\alpha$, it can be shown that this reduces the computational effort from $O(N^{2})$ to $O(N^{3/2})$ per step. Note that for any choice of $\alpha$ above formulas are correct; hence, the choice of $\alpha$ only influences which cutoff $r_{c}$ and $k_{c}$ can be chosen for a desired accuracy. In practice, $\alpha$ is by default heuristically determined from the prescribed tolerance $\delta$ on the energy approximation and a fixed cutoff $r_{c}$ (typically the one used for the van der Waals interactions) as[3]
@@
To further speed-up computation, it is possible to carry out the reciprocal sum using fast Fourier transform (FFT) by distributing the charges onto a lattice via spline interpolation. This significantly speeds up the calculation of $U_{\mathrm{recip.}}$, and the overall complexity reduces to $O(N\log N)$. Following OpenMM, we implement smooth particle-mesh Ewald (PME).[29, 3] For this, we also implement B-splines[30] based on the recursive definition of the splines, resulting in custom forward and backward implementation within PyTorch. The number of nodes in the mesh is chosen heuristically as $n_{\mathrm{mesh}}=2\alpha d/3\delta^{1/5}$, where $\alpha$ is again determined from Eq. (13).
Exclusions and Exceptions:
@@
Atoms that are directly bonded or indirectly bonded by one additional covalent bond to each other are excluded from the non-bonded energy terms. In addition, atoms connected via three bonds (such as in torsion contributions) are often treated differently. The most common way to handle these 11-44 exceptions is to scale the strength of the non-bonded interactions by some common factor <1<1. Alternatively, some parameterizations of force-fields scale each exclusion by an individual factor. In DIMOS, exclusions and exceptions are handled at the level of neighborlists, where excluded interaction pairs are filtered out. Since all possible 11-44 exceptions are known at simulation setup, they are handled explicitly by the non-bonded interaction method.
II.2 Machine Learning Based Descriptions of Intermolecular Interactions
@@
MLIPs do typically not introduce an explicit distinction between non-bonded and bonded atoms, and instead often build higher-order interactions between all atoms within a cutoff distance using various approaches, see Ref. 19 and references therein. Usually, these approaches include physical features in some capacity, for example a large body of work only considers features that are invariant to rotations and translations of the system and uses rather standard ML architectures with learnable parameters.[31] Another line of work explicitly constructs the architectures such that their internal features or their output transform equivariantly to the input, i.e., for example rotating the system also leads to a predictable rotation of the output.[6, 32, 33]
@@
We implement the interface to two current state-of-the-art MLIP models: MACE[6] and ORB.[32] We want, however, to stress that DIMOS makes no explicit assumption on the model and is rather agnostic to the details, as long as the models are implemented in PyTorch. In Fig. 1(b) we schematically demonstrate the basic setup of message-passing graph neural network MLIPs, in which many-body interaction terms are built using successive message passing iterations. However, since we do not implement any novel MLIP architecture in DIMOS, we refer to the literature on details as well as advantages/disadvantages of the respective approaches.[34, 19]
@@
Using DIMOS, the best performance is achieved if the model expects a combination of the neighborlist, distance vectors or distances, and/or shifts, since they are calculated anyway in DIMOS and recalculating (some of them) within the model leads to wasted calculations. By having the full simulation framework implemented in PyTorch, it is possible to compile parts or the whole simulation. This also extends to compiling MLIPs using torch.compile, which is currently not possible when using LAMMPS[2] or OpenMM[3] to interface the MLIP. Instead, these simulation engines require the use of TorchScript, forcing users to rewrite model code to conform to a restricted subset of Python in order to convert the model into a static graph. Albeit this does not necessarily impose a performance penalty, it makes the code more rigid and often harder to debug, due to loosing many of Python’s dynamic features.
Setting up a system modeled by a (foundational) MACE model can be achieved in a single line:
@@
In addition, there are some optional parameters, which, however, are not necessary in general. This highlights the simplicity of using DIMOS to run simulations using an MLIP, compared to other established approaches, often requiring a larger number of intermediate steps. For example, using the model in LAMMPS[2] requires implementing it using TorchScript in the first place (as discussed above) and compiling LAMMPS from scratch using either the ML-IAP extension or the specific branch maintained by the MACE developers, then exporting the model in a specific way, generating a corresponding input file, and transforming the coordinate file into a usable format.
@@
MACE is currently one of the most prominent MLIPs in the literature,[6] with diverse applications in different domains. We specifically employ the MACE material project foundation model (mace-mp),[35] which is available in the small, medium, and large variant depending on the chosen parameterization. In DIMOS, we support the acceleration of MACE using NVIDIA cuEquivariance, [^2] See https://github.com/NVIDIA/cuEquivariance for more details. and activate it in all our benchmarks.
II.3 Machine Learning/Molecular Mechanics